News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 12 April 2002 9.00am EST

During the last week the observing schedule was replanned twice to include two observations accepted as Targets Of Opportunity (TOO). The Gamma-ray Burster GRB020405 was accepted on 5 April and observed on 6 April, and the X-ray transient XTE J1751-305, also accepted on 5 April was observed on 10 April. Observations of the objects impacted by the TOO's will be observed in a later schedule.

A real-time procedures was completed on 10 April to update the on-board Ephemeris.

A press release was made on March 10 in conjunction with a NASA Space Science Update announcing observations of two neutron stars that suggest that the matter in the stars is denser that the standard models predict. One possibility is that the matter is in the form of strange quark matter. The release generated a great deal of interest in the press, and can be found at:

The observing schedule for the next week is shown below and includes two constrained observations of the comet C/2002 C1.

C/2002 C1               ACIS-S       Apr 14 
XTE J1908+094           ACIS-S       Apr 15 
Radiation Belts
C/2002 C1               ACIS-S
Field-142549+353248     ACIS-I
Radiation Belts                      Apr 17 
NGC 4038/NGC 4039       ACIS-S       Apr 18 
SGR A*                  ACIS-I       Apr 19 
1343-601                ACIS-S
0858-771                ACIS-S
1711-339                ACIS-S
M 2-9                   ACIS-S
BR1600+0724             ACIS-S
Tornado                 ACIS-I
RX J170930.2-263927     ACIS-S       Apr 20 
PSS 1605+4405           ACIS-S
Radiation Belts
Capella (2 obs)         HRC-S/LETG
Field-142549+35324      ACIS-I       Apr 21


All spacecraft subsystems continued to operate nominally.

Return to Status Reports